# Continuous Diffusion for Mixed-Type Tabular Data

### v2

@@ 4.2 Results
Table 1 shows the ranks of all generative averaged across all datasets for all considered metrics. The ranks in terms of the F1 and AUC scores are averaged over the classification task datasets. Likewise, the RMSE rank averages include the regression task datasets only. We assign the maximum possible rank when a model was not trainable on a given dataset. This includes TabDDPM, which outputs NaNs for acsincome and diabetes and CoDi, which we consider to be prohibitively expensive to train on diabetes (estimated 14.5 hours) and lending (estimated 60 hours). The dataset-specific results (including standard errors) and average metrics over all datasets are detailed in Appendix N. We provide visualizations of the captured correlations in the synthetic sample compared to the real training set in Appendix M and distribution plots for a qualitative comparison in Appendix L.
@@
CDTD consistently outperforms the considered benchmark models in most sample quality metrics. We see a major performance edge in terms of the detection score, the L2 distance of the correlation matrices and the regression-based metrics. Only for the Jensen-Shannon divergence, CDTD is dominated by ARF, a tree-based method that is expected to model categorical features particularly well. Interestingly, CDTD performs similar to TabDDPM on F1 scores but outperforms it dramatically for regression tasks. TabDDPM appears to favor modeling categorical features accurately, thereby sacrificing continuous features, as visualized in Appendix M. By utilizing score interpolation, CDTD is able to model intricate correlation structure more accurate than other (multinomial diffusion-based) frameworks. Most importantly, type-specific noise schedules always outperform the feature-specific noise schedule variant and often also the single noise schedule version. This illustrates the importance of accounting for the high heterogeneity in tabular data on the feature type level. Imposing different noise schedules per feature, however, appears to force too many constrains on the model and hence decreases sample quality. Per-feature noise schedules would requires more training steps to converge, as can be seen in Appendix K.
Table 1: Average performance rank of each generative model across eleven datasets. Per performance metric, bold indicates the best, underline the second best result. CoDi is prohibitively expensive to train on lending and diabetes, TabDDPM produces NaNs for acsincome and diabetes, and therefore get assigned rank 8 for these datasets.
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
RMSE $3.000_{\pm 1.673}$ $6.400_{\pm 1.356}$ $6.200_{\pm 1.833}$ $6.400_{\pm 1.200}$ $5.600_{\pm 1.356}$ $\mathbf{2.400}_{\pm 0.490}$ $\underline{2.600}_{\pm 1.855}$ $3.400_{\pm 2.059}$
F1 $5.167_{\pm 1.951}$ $6.667_{\pm 1.374}$ $6.667_{\pm 0.745}$ $\underline{3.333}_{\pm 2.285}$ $5.333_{\pm 2.494}$ $\underline{3.333}_{\pm 1.247}$ $\mathbf{2.667}_{\pm 1.106}$ $3.000_{\pm 1.915}$
AUC $4.500_{\pm 1.607}$ $7.000_{\pm 1.155}$ $6.500_{\pm 0.764}$ $4.167_{\pm 2.115}$ $6.000_{\pm 2.160}$ $\underline{2.333}_{\pm 1.599}$ $\mathbf{2.167}_{\pm 0.898}$ $3.500_{\pm 1.500}$
L2 dist. of corr. $4.455_{\pm 1.305}$ $6.727_{\pm 1.213}$ $6.455_{\pm 1.157}$ $5.636_{\pm 2.346}$ $5.545_{\pm 1.876}$ $\underline{2.545}_{\pm 1.305}$ $\mathbf{2.091}_{\pm 0.793}$ $2.636_{\pm 1.553}$
Detection score $4.818_{\pm 1.336}$ $7.182_{\pm 1.113}$ $5.636_{\pm 0.881}$ $4.000_{\pm 2.412}$ $6.455_{\pm 1.777}$ $3.273_{\pm 0.862}$ $\mathbf{1.818}_{\pm 1.696}$ $\underline{2.909}_{\pm 1.505}$
JSD $\mathbf{1.273}_{\pm 0.617}$ $6.818_{\pm 1.113}$ $7.182_{\pm 0.716}$ $5.818_{\pm 1.641}$ $5.909_{\pm 1.083}$ $\underline{2.364}_{\pm 0.979}$ $3.000_{\pm 0.603}$ $3.727_{\pm 1.052}$
WD $4.273_{\pm 1.483}$ $6.000_{\pm 1.758}$ $6.455_{\pm 1.157}$ $5.000_{\pm 2.449}$ $6.727_{\pm 1.286}$ $3.000_{\pm 1.348}$ $\mathbf{2.273}_{\pm 1.052}$ $\underline{2.364}_{\pm 1.494}$
DCR $4.909_{\pm 1.564}$ $6.818_{\pm 1.466}$ $5.182_{\pm 2.443}$ $3.636_{\pm 2.837}$ $5.364_{\pm 1.920}$ $3.727_{\pm 2.049}$ $\underline{3.364}_{\pm 1.611}$ $\mathbf{3.091}_{\pm 1.443}$
Figure 3: Average wall-clock time for training and for sampling 1000 samples (excluding acsincome, diabetes and lending).
Training and sampling time.
Figure 3 shows the average wall-clock time over all datasets (for which training was feasible for all models) for training as well as the time for sampling 1000 data points of each baseline model and the most complex CDTD variant. Details are given in Appendix P. The use of embeddings (instead of one-hot encoding) for categorical features drastically reduces the training time. CDTD trains the fastest on average and is able to reach a maximum training time of less than 30 min across all datasets. Thus, CDTD exhibits improved scaling to an increase in the number of categories. The ODE definition of the diffusion process also implies competitive sampling speeds compared to TVAE and CTGAN and in particular, to the diffusion-based benchmarks CoDi and TabDDPM. In light of the vastly improved scalability, the CDTD results in Table 1 could in practice trivially be improved further by simply scaling up the architecture.
@@
We conduct an ablation study over four datasets to investigate the separate components of our CDTD framework. The results are given in Table 2 (detailed results are in Appendix O). The baseline model Config. A includes a single noise schedule with the original piece-wise linear formulation (Dieleman et al., 2022) without loss normalization, improved model initialization or adaptive normalization, and the CE and MSE losses are naively averaged. Note that this configuration still is a novel contribution to the literature. Config. B adds our feature homogenization (i.e., loss normalization, improved initialization and adaptive normalization schemes), Config. C adds our proposed functional form for a single noise schedule with uniform initialization, and Config. D imposes per-type noise schedules. Lastly, we add the suggested (low noise level) overweighting timewarping initialization to arrive at the full CDTD (per type) model. We see the switch from the piece-wise linear functional form to our more robust noise schedule variant slightly harms sample quality. However, the per-type variant and the more informed initialization scheme compensate for this difference. Main differences are in the RMSE and detection score as well as training efficiency (the loss calibration and improved initialization facilitate model convergence). The final model especially works well on the larger datasets compared to the baseline (see Appendix O), as smaller datasets are relatively easy to fit with 3 million parameters, even without any model improvements.
@@ 5 Conclusion and Discussion
We propose a Continuous Diffusion model for mixed-type Tabular Data (CDTD) that combines score matching and score interpolation and imposes Gaussian diffusion processes on both continuous and embedded categorical features. We compared to various benchmark models and to a single noise schedule as typically used in image diffusion models. Our results indicate that acknowledging the high feature heterogeneity in tabular data on the feature type level and aligning type-specific diffusion elements, such as the noise schedules or losses, substantially benefits sample quality. CDTD also shows vastly improved scalability and can accommodate an arbitrary number of categories.
Our paper serves as an important step to adopting the diffusion probabilistic framework customized to tabular data. In particular, the common type of noise schedules allows for a framework that is easy to extend and may accelerate progress on diffusion models for tabular data. Crucially, CDTD allows the direct application of diffusion-related advances from the image domain, like classifier-free guidance, to tabular data. We leave further extensions, e.g., the exploration of accelerated sampling, efficient score model architectures, or different forms of adaptive noise schedules, to future work.
Finally, we want to highlight that synthetic data could be misused to support unwarranted claims. In any case, any generated data should not be blindly trusted and synthetic data based inferences should be compared to results derived from the real data. On the other hand, the correct usage of generative models can allow for better privacy preservation, facilitate data sharing and open science practices.
@@
The main limitation of the CDTD model is the addition of another hyperparameter, i.e., the type of noise schedule, and tuning a hyperparameter of a generative model can be a costly endeavor. However, our results also show that a per type schedule is most often optimal. Dieleman et al. (2022) show that the results of score interpolation can be sensitive to the initialization of the embeddings. We have not encountered any problems with the same standard deviation of 0.001 for all datasets. However, this may not apply in general. While the DCR indicates no privacy issues for the benchmark datasets used, additional caution must be taken when generating synthetic data from privacy sensitive sources. Lastly, our model may be outperformed by other generative models specialized for different types of tabular data, such as time-series.
@@ Appendix A Loss Calibration
A priori, we let the model be indifferent between features. Therefore, we aim to scale the loss of each feature such that they attain the same loss at the terminal timestep $t=1$. In that case, the signal-to-noise ratio is sufficiently low to approximate a situation in which the model has no information about the data. We are looking for calibrated losses $\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)$ and $\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},1)$ which in expectation achieve a unit loss at $t=1$. This way, the model treats all features equally before having seen any information.
For a single scalar feature and given timestep $t$, we can write the empirical denoising score matching loss (see Equation 3) in the EDM parameterization (Karras et al., 2022) as:
@@
where $F_{\bm{\theta}}^{(i)}$ denotes the neural network output for feature $i$ that parameterizes the score model $s_{\bm{\theta}}$. The parameters $c_{\text{skip}}(t)=\sigma_{\text{data}}^{2}/(\sigma^{2}(t)+\sigma_{\text{data}}^{2})$ and $c_{\text{out}}(t)=\sigma(t)\cdot\sigma_{\text{data}}/(\sqrt{\sigma^{2}(t)+\sigma_{\text{data}}^{2}})$ depend on $\sigma(t)$ (and $\sigma_{\text{data}}$) and therefore on timestep $t$. For $t\rightarrow 1$, $\sigma(t)$ approaches the maximum noise level $\sigma_{\text{cont},\max}$ and $c_{\text{skip}}(t)\rightarrow 0$ and $c_{\text{out}}(t)\rightarrow 1$ such that the score model directly predicts the data at high noise levels. For $t\rightarrow 0$, the model shifts increasingly towards predicting the error that has been added to the true data. In the EDM parameterization, the explicit timestep weight (used to achieve a unit loss across timesteps at initialization (see Appendix B) is $\lambda(t)=1/c_{\text{out}}(t)^{2}\approx 1$ for $t=1$.
@@ Appendix B Output Layer Initialization
We initialize the neural network such that $F_{{\bm{\theta}},c}^{(j)}=\log p_{c}$ for all $c$. This can be achieved by initializing the output layer weights to zero and the output layer biases to the relevant training set log-proportions of the corresponding class. Hence, this initialization gives us
@@ Appendix C Adaptive Normalization of the Average Diffusion Loss
Both the loss calibration (see Appendix A) and output layer initialization (see Appendix B) ensure that the losses across timesteps (and features) are equal at initialization. During training, the adaptive noise schedules allow the model to focus automatically on the noise levels that matter most, i.e., where the loss increase is steepest. However, the better the model becomes at a given timestep $t$, the lower the loss at that timestep, and the lower the gradient signal relative to the signal received for timesteps $\tilde{t}>t$. To counteract this, we additionally use an adaptive normalization of the average diffusion loss (averaged over the features) across timesteps. Specifically, we want to weight the average diffusion loss at timestep $t$, $\mathcal{L}(t)$ given in Equation 6, such that the normalized loss is the same (equal to one) for all $t$. Similar methods have been used by Karras et al. (2023) and Kingma & Gao (2023), we follow the latter in the setup of the corresponding network.
We train a neural network alongside our diffusion model to predict $\mathcal{L}(t)$ based on $t$ and using the MSE loss to learn this weighting. In particular, we first compute $c_{\text{noise}}(t)=\log(t)/4$ following the EDM parameterization (Karras et al., 2022). Then, we embed $c_{\text{noise}}$ in frequency space (1024-dimensional) using Fourier features. The result is then passed through a single linear layer to output a scalar value, which is in turn passed through an exponential function to ensure that the prediction ^$\hat{\mathcal{L}}(t)\geq 0$. We initialize the weights and biases to zero, to ensure that at model initialization we have a unit normalization.

### v3

@@ 4.2 Results
Table 1: Average performance rank of each generative model across eleven datasets. Per metric, bold indicates the best, underline the second best result. We assigned the rank 10 for CoDi on lending and diabetes, TabDDPM on acsincome and diabetes, SMOTE on acsincome and covertype.
SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
RMSE $\underline{3.400}_{\pm 3.382}$ $3.800_{\pm 2.482}$ $7.800_{\pm 1.470}$ $7.800_{\pm 2.227}$ $8.200_{\pm 1.470}$ $6.800_{\pm 1.939}$ $6.800_{\pm 1.720}$ $\mathbf{3.000}_{\pm 0.632}$ $\underline{3.400}_{\pm 2.059}$ $4.200_{\pm 2.227}$
F1 $3.667_{\pm 3.145}$ $6.333_{\pm 2.285}$ $8.333_{\pm 1.491}$ $8.000_{\pm 1.000}$ $4.167_{\pm 2.794}$ $6.667_{\pm 3.091}$ $7.500_{\pm 1.607}$ $3.833_{\pm 1.462}$ $\mathbf{3.167}_{\pm 1.344}$ $\underline{3.500}_{\pm 1.979}$
AUC $4.667_{\pm 2.749}$ $5.667_{\pm 2.055}$ $8.667_{\pm 1.106}$ $7.833_{\pm 1.067}$ $4.833_{\pm 2.794}$ $7.500_{\pm 2.872}$ $7.333_{\pm 1.374}$ $\underline{2.500}_{\pm 1.500}$ $\mathbf{2.333}_{\pm 0.943}$ $3.833_{\pm 1.675}$
L2 dist. of corr. $4.818_{\pm 2.918}$ $5.636_{\pm 1.872}$ $8.091_{\pm 1.781}$ $7.909_{\pm 1.564}$ $7.000_{\pm 3.191}$ $6.909_{\pm 2.429}$ $6.818_{\pm 1.402}$ $\underline{2.727}_{\pm 1.286}$ $\mathbf{2.273}_{\pm 0.862}$ $3.000_{\pm 1.595}$
Detection score $3.909_{\pm 3.502}$ $6.182_{\pm 1.696}$ $8.818_{\pm 1.466}$ $7.273_{\pm 1.213}$ $5.000_{\pm 3.045}$ $8.091_{\pm 2.391}$ $6.000_{\pm 1.595}$ $3.909_{\pm 1.164}$ $\mathbf{2.455}_{\pm 1.827}$ $\underline{3.545}_{\pm 1.725}$
JSD $7.182_{\pm 2.167}$ $\mathbf{1.273}_{\pm 0.617}$ $8.182_{\pm 1.641}$ $8.818_{\pm 1.029}$ $6.909_{\pm 2.314}$ $7.000_{\pm 1.651}$ $6.545_{\pm 1.305}$ $\underline{2.455}_{\pm 1.076}$ $3.091_{\pm 0.793}$ $3.727_{\pm 1.052}$
WD $\underline{3.091}_{\pm 3.315}$ $5.636_{\pm 1.611}$ $7.545_{\pm 1.827}$ $8.000_{\pm 1.477}$ $6.455_{\pm 3.144}$ $8.364_{\pm 1.823}$ $5.727_{\pm 2.339}$ $4.182_{\pm 1.466}$ $3.182_{\pm 1.192}$ $\mathbf{3.000}_{\pm 1.954}$
DCR $6.000_{\pm 2.558}$ $6.182_{\pm 2.328}$ $8.455_{\pm 1.725}$ $6.182_{\pm 3.186}$ $4.455_{\pm 3.726}$ $6.545_{\pm 2.426}$ $5.909_{\pm 1.676}$ $4.091_{\pm 2.678}$ $\underline{3.818}_{\pm 2.124}$ $\mathbf{3.545}_{\pm 1.924}$
Table 1 shows the average rank of each generative model across all datasets for the considered metrics. The ranks in terms of the F1 and AUC scores are averaged over the classification task datasets. Likewise, the RMSE rank averages include the regression task datasets. We assign the maximum possible rank when a model could not be trained on a given dataset or could not be evaluated in reasonable time. This includes TabDDPM, which outputs NaNs for acsincome and diabetes and CoDi, which we consider to be prohibitively expensive to train on diabetes (estimated 14.5 hours) and lending (estimated 60 hours). Similarly, SMOTE is very inefficient in sampling for large datasets (78 min for 1000 samples on acsincome and 182 min on covertype) and does not finish the evaluation within 12 hours. The dataset-specific results (including standard errors) and average metrics over all datasets are detailed in Appendix R. We provide visualizations of the captured correlations in the synthetic sample compared to the real training set in Appendix Q and distribution plots for a qualitative comparison in Appendix P.
@@
CDTD consistently outperforms the considered benchmark models in most sample quality metrics. Specifically, we see a major performance edge in terms of the detection score, the L2 distance of the correlation matrices and the regression-based metrics. Only for the Jensen-Shannon divergence ARF, a tree-based method that is expected to model categorical features particularly well, outperforms CDTD. Interestingly, CDTD performs similar to TabDDPM on F1 scores, but outperforms it dramatically for regression tasks. TabDDPM appears to favor modeling categorical features accurately, thereby sacrificing continuous features, as visualized in Appendix Q. TabSyn, a latent-space diffusion model, performs worse than CDTD and often TabDDPM, which define diffusion in data space. In Appendix M, we further compare CDTD and TabSyn and investigate the benefits of defining a diffusion model in data space. By utilizing score interpolation, CDTD is able to model intricate correlation structure more accurately than other frameworks. Most importantly, type-specific noise schedules mostly outperform the feature-specific and single noise schedule variants. This illustrates the importance of accounting for the high heterogeneity in tabular data on the feature type level. The different noise schedules per feature, however, appear to force too many constraints on the model and thus, decrease sample quality. Per-feature noise schedules would require more training steps to converge, as can be seen in Appendix O.
@@
Figure 3: Average training and sampling wall-clock time for 1000 samples (excl. acsincome, diabetes, lending).
Training and sampling time.
Figure 3 shows the average wall-clock time over all (for all models feasible) datasets for training as well as the time for sampling 1000 data points for each baseline model and the per feature CDTD variant (see Appendix T for details). We exclude SMOTE due to its considerably longer sampling with an average of 1377 seconds for 1000 samples. CDTD’s use of embeddings (instead of one-hot encoding) for categorical features drastically reduces training times and thus, improves scaling to increasing number of categories. The ODE formulation of the diffusion process implies competitive sampling speeds, in particular compared to the diffusion-based benchmarks CoDi, TabDDPM and TabSyn. Despite TabSyn utilizing a separately trained encoder, this does not result in a lower-dimensional latent space and therefore, does not speed up sampling.
@@
We conduct an ablation study over four datasets to investigate the separate components of our CDTD framework. The results are given in Table 2 (detailed results are in Appendix S). The baseline model Config. A includes a single noise schedule with the original piece-wise linear formulation (Dieleman et al., 2022) without loss normalization, improved model initialization or adaptive normalization, and the CE and MSE losses are naively averaged. Note that this configuration still is a novel contribution to the literature. Config. B adds our feature homogenization (i.e., loss normalization, improved initialization and adaptive normalization schemes), Config. C adds our proposed functional form for a single noise schedule with uniform initialization, and Config. D imposes per-type noise schedules. Lastly, we add the suggested (low noise level) overweighting timewarping initialization to arrive at the full CDTD (per type) model. We see the switch from the piece-wise linear functional form to our more robust noise schedule variant slightly harms sample quality. However, the per-type variant and the more informed initialization scheme compensate for this difference. Main differences are in the RMSE and detection score as well as training efficiency (the loss calibration and improved initialization facilitate model convergence). The final model especially works well on the larger datasets compared to the baseline (see Appendix S), as smaller datasets are relatively easy to fit with 3 million parameters, even without any model improvements. We investigate the sensitivity of CDTD to important hyperparameters in Appendix L.
@@ 5 Conclusion and Discussion
We propose a Continuous Diffusion model for mixed-type Tabular Data (CDTD) that combines score matching and score interpolation and imposes Gaussian diffusion processes on both continuous and embedded categorical features. We compared CDTD to various benchmark models and to a single noise schedule as typically used in image diffusion models. Our results indicate that addressing the high feature heterogeneity in tabular data on the feature type level and aligning type-specific diffusion elements, such as the noise schedules or losses, substantially benefits sample quality. Moreover, CDTD shows vastly improved scalability and can accommodate an arbitrary number of categories.
Our paper serves as an important step to customizing the diffusion probabilistic framework to tabular data. In particular, the common type of noise schedules allows for an easy to extend framework that might accelerate progress on diffusion models for tabular data. Crucially, CDTD allows the direct application of diffusion-related advances from the image domain, like classifier-free guidance, to tabular data without the need for a latent encoding. We leave further extensions to the tabular data domain, e.g., the exploration of accelerated sampling, efficient score model architectures, different forms of adaptive noise schedules, or the adaption to the data imputation task for future work.
Finally, we want to emphasize the potential misuse of synthetic data to support unwarranted claims. Any generated data should therefore not be blindly trusted, and synthetic data based inferences should always be compared to results from the real data. However, the correct use of generative models enables better privacy preservation and facilitates data sharing and open science practices.
@@
The main limitation of CDTD is the addition of hyperparameters, and tuning hyperparameters of a generative model can be a costly endeavor. However, our results also show that (1) a per type schedule is most often optimal and (2) our default hyperparameters perform well across a diverse set of datasets. Dieleman et al. (2022) show that the results of score interpolation for text data can be sensitive to the initialization of the embeddings. We have not encountered similar problems on tabular datasets (see Table 7). While the DCR indicates no privacy issues for the benchmark datasets used, additional caution must be taken when generating synthetic data from privacy sensitive sources. Lastly, for specific types of tabular data, such as time-series, our model may be outperformed by other generative models specialized for that type. While CDTD could be directly used for imputation using RePaint (Lugmayr et al., 2022), a separate training process is required to achieve the best results (Liu et al., 2024). Therefore, we leave the adaption of CDTD to the imputation task for future work.
@@ Appendix A Loss Calibration
A priori, we let the model be indifferent between features, that is, we scale the loss of each feature such that at the terminal timestep the same loss is attained. Here, the signal-to-noise ratio is sufficiently low to approximate a situation in which the model has no information about the data. Thus, we are looking for calibrated losses $\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)$ and $\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},1)$ which at $t=1$ achieve unit loss in expectation.
For a single scalar feature and a given timestep $t$, we can write the empirical denoising score matching loss (Equation 3) in the EDM parameterization (Karras et al., 2022) as:
@@
where $F_{\bm{\theta}}^{(i)}$ denotes the neural network output for feature $i$ that parameterizes the score model $s_{\bm{\theta}}$. The parameters $c_{\text{skip}}(t)=\sigma_{\text{data}}^{2}/(\sigma^{2}(t)+\sigma_{\text{data}}^{2})$ and $c_{\text{out}}(t)=\sigma(t)\cdot\sigma_{\text{data}}/(\sqrt{\sigma^{2}(t)+\sigma_{\text{data}}^{2}})$ depend on $\sigma(t)$ (and $\sigma_{\text{data}}$) and therefore on timestep $t$. For $t\rightarrow 1$, $\sigma(t)$ approaches the maximum noise level $\sigma_{\text{cont},\max}$ and $c_{\text{skip}}(t)\rightarrow 0$ and $c_{\text{out}}(t)\rightarrow 1$ such that the score model directly predicts the data at high noise levels. For $t\rightarrow 0$, the model shifts increasingly towards predicting the error that has been added to the true data. In the EDM parameterization, the explicit timestep weight (used to achieve a unit loss across timesteps at initialization, see Appendix B) is $\lambda(t)=1/c_{\text{out}}(t)^{2}\approx 1$ for $t=1$.
@@ Appendix B Output Layer Initialization
We initialize the neural network such that $F_{{\bm{\theta}},c}^{(j)}=\log p_{c}$ for all $c$. This is achieved by initializing the output layer weights to zero and the output layer biases to the relevant training set log-proportions of the corresponding class. Hence, this initialization gives us
@@ Appendix C Adaptive Normalization of the Average Diffusion Loss
Both the loss calibration (see Appendix A) and output layer initialization (see Appendix B) ensure that the losses across timesteps (and features) are equal at initialization. During training, the adaptive noise schedules allow the model to focus automatically on the noise levels that matter most, i.e., where the loss increase is steepest. However, the better the model becomes at a given timestep $t$, the lower the loss at the respective timestep, and the lower the gradient signal relative to the signal for timesteps $\tilde{t}>t$. We counteract this with adaptive normalization of the average diffusion loss (averaged over the features) across timesteps. Specifically, we want to weight the average diffusion loss at timestep $t$, $\mathcal{L}(t)$ given in Equation 6, such that the normalized loss is the same (equal to one) for all $t$. Similar methods have been used by Karras et al. (2023) and Kingma & Gao (2023), we follow the latter in the setup of the corresponding network.
We train a neural network alongside our diffusion model to predict $\mathcal{L}(t)$ based on $t$ and use the MSE loss to learn this weighting. First, we compute $c_{\text{noise}}(t)=\log(t)/4$ following the EDM parameterization (Karras et al., 2022). Then, we embed $c_{\text{noise}}$ in frequency space (1024-dimensional) using Fourier features. The result is passed through a single linear layer to output a scalar value, passed through an exponential function to ensure that the prediction ^$\hat{\mathcal{L}}(t)\geq 0$. We initialize the weights and biases to zero, to ensure that at model initialization we have a unit normalization.